### v1

# IEEEexample:BSTcontrol

Abstract
We study parameterizations of stabilizing nonlinear policies for learning-based control. We propose a structure based on a nonlinear version of the Youla-Kučera parameterization combined with robust neural networks such as the recurrent equilibrium network (REN). The resulting parameterizations are unconstrained, and hence can be searched over with first-order optimization methods, while always ensuring closed-loop stability by construction. We study the combination of (a) nonlinear dynamics, (b) partial observation, and (c) incremental closed-loop stability requirements (contraction and Lipschitzness). We find that with any two of these three difficulties, a contracting and Lipschitz Youla parameter always leads to contracting and Lipschitz closed loops. However, if all three hold, then incremental stability can be lost with exogenous disturbances. Instead, a weaker condition is maintained, which we call d-tube contraction and Lipschitzness. We further obtain converse results showing that the proposed parameterization covers all contracting and Lipschitz closed loops for certain classes of nonlinear systems. Numerical experiments illustrate the utility of our parameterization when learning controllers with built-in stability certificates for: i) “economic” rewards without stabilizing effects; ii) short training horizons; and iii) uncertain systems.
{IEEEkeywords}
Reinforcement learning, Youla parameterization, recurrent equilibrium networks, contraction, IQC
1 Introduction
Deep reinforcement learning (RL) is an emerging technology for general-purpose nonlinear control design via simulation. It has been successfully applied in many complex domains ranging from strategy games [1] to robotics [2, 3] to nuclear fusion [4]. The standard approach is to minimize empirical estimates of an expected cost over repeated episodes with random disturbances, typically using deep neural networks for black-box policy parameterizations [5]. A key factor enabling the broad application and scalability of deep RL is the use of simple, gradient-based optimization methods to train policies [6]. However, black-box approaches such as deep RL are limited by a lack of guarantees on the stability, response to disturbances, and sensitivity to model error of the closed-loop system. While the risk of instability can be reduced by using long training horizons with regulation or tracking-style costs, closed-loop guarantees are not explicitly part of the standard RL problem formulation, which can lead to unexpected failure modes [7, 8]. We seek to remedy this limitation while retaining the benefits of simple gradient-based RL tools — i.e., without additional computationally-expensive constraints or stability analysis procedures during training.
The purpose of this paper is to construct policy parameterizations that are stable by design — i.e., closed-loop stability and robustness constraints are embedded directly within the policy architecture itself. Learning controllers from a class of stable-by-design policies decouples closed-loop stability from the optimization algorithms, reward functions, and data used to train the policy, enabling the learning of high-performance controllers without risking instability.
We construct a stable-by-design policy parameterization based on the classical Youla-Kučera parameterization from linear systems theory [9, 10]. We study nonlinear versions of the Youla parameterization and present novel results extending it to strong, incremental notions of stability and robustness which are suitable for machine learning applications. We then provide a constructive realization of the parameterization called the Youla-REN, which fuses the Youla framework with the recurrent equilibrium network (REN) class of robust neural models [11]. Combined with the built-in behavioral properties of RENs, all Youla-REN policies are naturally closed-loop stabilizing and robust, making them suitable for plug-and-play learning with data-driven control frameworks such as deep RL.
(a) RL.
(b) Residual RL.
(c) Youla-REN.
Figure 1: Feedback controller architectures for deep RL.
1.1 Feedback Policy Parameterizations
Fig. 1(a) shows the typical parameterization of an RL policy as a neural network in feedback with a (possibly nonlinear, unstable) system. With black-box policies, this architecture offers no closed-loop stability or robustness guarantees, and does not incorporate prior knowledge of the system.
One way to take advantage of prior knowledge is to design a “base controller” — for example, using classical model-based techniques — and augment it with a neural network as in Fig. 1(b). This is called residual RL [12, 13, 14]. The base controller is used to bootstrap performance and allows us to naturally fuse model-based control with data-driven RL. However, closed-loop stability and robustness constraints are typically not imposed in residual RL. It would be possible to do so via (for example) the small gain theorem if a gain bound from the neural network output to the system output is known and gain-bounded neural networks are employed. However, this will typically be very conservative.
We can reduce conservatism by adding a third block to the parameterization — an observer — as shown in Fig. 1(c), which requires further prior knowledge in the form of a system model. Observers can be thought of as “model-pass filters.” They let through information that is consistent with their internal model and filter out unexpected information from the measurements (a.k.a the innovations or “surprises”). We say that a neural network reacts to surprises if we feed it differences between the measurements and observer predictions. This is a version of the Youla parameterization, and the neural network in Fig. 1(c) is called the Youla parameter. Its main attraction is that, for certain classes of systems and observers: (1) it automatically guarantees closed-loop stability without restricting the network gain, provided that the Youla parameter is itself stable; and (2) it is not a restrictive parameterization since it covers all stabilizing controllers for the system. For systems and observers where this is not the case, we can still apply the small-gain theorem to ensure stability. In the Youla parameterization, the gain from the network output to the surprises is typically much smaller than to the measurements themselves, significantly reducing conservatism.
1.2 The Youla Parameterization for Nonlinear Systems
The Youla parameterization [9, 10] was first established to parameterize all stabilizing linear controllers for a given linear system, and has long played a central role in linear robust and adaptive control theory [15, 16], although to our knowledge the use of the Youla parameterization in RL was first proposed in [17]. A common construction is to augment an observer-based stabilizing base controller with a stable linear system $\mathcal{Q}$ called the Youla parameter, which reacts to surprises much like in Fig. 1(c) [18]. $\mathcal{Q}$ is then optimized to boost the original closed-loop system performance while always preserving stability with hard guarantees. We note that parameterizations essentially equivalent to the Youla parameterization have been proposed under many names over the years, including: model reference transformations and $\mathcal{Q}$-parameterizations [19]; internal model control (IMC) [20]; disturbance feedback in model predictive control (MPC) [21]; and “nature’s y” in online learning approaches to control [22].
Extensions of the Youla parameterization to nonlinear systems have been studied for many years [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Early work addressed the problem by expressing the parameterization in terms of coprime factors [25, 26], which are difficult to realize for general nonlinear systems. [27, 28, 29, 30] proposed using kernel representations to enable intuitive realizations of the nonlinear Youla parameterization since they can be translated to observer-based stabilizing controllers in state-space. State-space formulations of the problem were also studied directly in [31, 32].
Each of [23, 24, 25, 26, 27, 28, 29, 30, 31, 32] focused on parameterizing asymptotically or finite-gain stabilizing controllers about a particular equilibrium. However, there are many different notions of stability for nonlinear systems [33]. In RL and learning-based control, we argue that the stronger notion of incremental stability is more relevant, since it ensures stable behavior when generalizing to unseen data. In this paper, we therefore seek a formulation of the Youla parameterization that parameterizes incrementally stabilizing controllers in the sense of contraction [34] and incremental integral quadratic constraints (IQCs) [35].
1.3 Direct Parameterization of Stable Nonlinear Systems
A bottleneck in applying nonlinear versions of the Youla parameterization has been the construction of the stable nonlinear parameter $\mathcal{Q}$. Prior work in system identification has introduced convex parameterizations of contracting and Lipschitz nonlinear dynamic models based on polynomials [36, 37] and neural networks [38]. However, these parameterizations were expressed in the form of linear matrix inequalities via sum-of-squares and IQC constructions, which limited their scalability and required custom optimization methods [39, 38].
A key recent advance was the so-called direct parameterization of contracting and Lipschitz models in the form of RENs [11]. In a direct parameterization, the parameters are unconstrained and standard first-order optimization tools such as stochastic gradient descent can be applied, enabling straightforward integration with existing RL frameworks. Moreover, RENs are expressive: they universally approximate all contracting and Lipschitz nonlinear systems [40]. Alternative parameterizations of RENs were introduced in [41] to significantly improve their scalability to large neural models.
1.4 Contributions & Paper Outline
This paper studies the parameterization of stable-by-design feedback controllers for partially-observed nonlinear systems. Our main contributions are:
  • We introduce a novel version of the Youla parameterization which guarantees contracting and Lipschitz closed loops in a disturbance-free setting, and prove that the effects of disturbances are constrained to bounded tubes around the disturbance-free trajectories.
  • We show that a sufficient condition for contracting and Lipschitz closed loops is that the innovations are decoupled from the Youla augmentation. If the two are coupled, we can still restrict the Youla parameter’s input-output properties via IQCs to achieve closed-loop stability.
  • We provide partial converse results proving that the parameterization contains all contracting and Lipschitz closed-loops for certain classes of nonlinear systems.
  • Combined with RENs, we show via numerical experiments that our resulting Youla-REN policies can achieve strong performance on deep RL tasks while preserving closed-loop stability under disturbances and uncertainty.
Prior work by the authors introduced preliminary versions of the Youla-REN for linear systems [42, 40, 11] and partially-observed nonlinear systems without disturbances [43]. This paper studies the general case of partially-observed nonlinear systems with disturbances, examines the potential loss of contraction and Lipschitzness in this case, introduces the new notion of d-tube contraction and Lipschitzness, and proves novel converse results.
Concurrent work in [44, 45] also extends the Youla parameterization to the contraction framework, and includes similar statements to Theorems 5.1 and 7 in the present paper, variants of which also appeared in [43]. However, [44, 45] do not consider the effect of exogenous disturbances on the closed-loop system. Further concurrent work [46, 47, 48] studies the Youla parameterization from the perspective of input-output operators, and also utilizes RENs as the stable parameter. However, [46] only considers the state-feedback setting, and [47, 48] are limited to finite-gain stability for stable plants and make strong assumptions about measurable disturbances for the results on closed-loop contraction and Lipschitzness.
The remainder of this paper is organized as follows. Section 2 formulates the problem of constructing robustly-stabilizing policy parameterizations. Section 3 introduces a nonlinear version of the Youla parameterization. Our main theoretical results are presented in Sections 46. Section 7 studies the Youla-REN via numerical examples. Conclusions and proofs are provided in Section 8 and Appendices .1.6.
1.5 Notation
Let $\mathbb{N}$ and ($\mathbb{R}^{+}$) $\mathbb{R}$ be the set of natural and (positive) real numbers, respectively. We denote the set of sequences $x:\mathbb{N}\rightarrow\mathbb{R}^{n}$ by $\ell^{n}$, where the superscript $n$ is omitted if it is clear from context. Let $\ell_{2}^{n}\subset\ell^{n}$ be the set of sequences with finite $\ell_{2}$ norm, i.e., $\|x\|:=\left(\sum_{t=0}^{\infty}|x_{t}|^{2}\right)^{\frac{1}{2}}<\infty$ for any $x\in\ell_{2}^{n}$ where ||$\cdot$|| ⋅ | is the Euclidean norm. Let $x_{i:j}:=(x_{i},x_{i+1},\ldots,x_{j})$ be the sequence truncation over $[i,j]$ with $i\leq j$. We use $\|x\|_{T}$ to denote the $\ell_{2}$ norm of $x_{0:T}$ with $T\in\mathbb{N}$. We write $A(\succeq)\succ 0$ for positive (semi-)definite matrices. A function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is said to be Lipschitz continuous if $\exists\,\gamma\in\mathbb{R}^{+}$ such that $|f(a)-f(b)|\leq\gamma|a-b|$ for all $a,b\in\mathbb{R}^{n}$. We denote the concatenation of vectors $u\in\mathbb{R}^{n}$, $v\in\mathbb{R}^{m}$ as $[u;v]$.
2 Problem Formulation
2.1 Problem Statement
We consider nonlinear systems $\mathcal{G}$ of the form
$\mathcal{G}:\quad\left\{\begin{aligned} x_{t+1}&=f(x_{t},\eta_{t},u_{t})+w_{t}\\ y_{t}&=h(x_{t})+v_{t}\end{aligned}\right.$ (1)
with internal states $x_{t}\in\mathbb{R}^{n_{x}}$, controlled inputs $u_{t}\in\mathbb{R}^{n_{u}}$, and measured outputs $y_{t}\in\mathbb{R}^{n_{y}}$. The states and measurements are perturbed by (unknown) additive process disturbances $w_{t}\in\mathbb{R}^{n_{x}}$ and measurement noise $v_{t}\in\mathbb{R}^{n_{y}}$, respectively. Here $\eta_{t}\in\mathbb{R}^{n_{\eta}}$ are known external inputs representing reference signals, feedforward commands, disturbance previews, or other known inputs. We denote the unknown disturbances $d_{t}=[w_{t};v_{t}]$ and the controlled outputs $z_{t}=[x_{t};u_{t}]$.
Our task is to parameterize stabilizing feedback policies $u=\mathcal{K}(\eta,y)$ for (1). Specifically, our goal is to construct a policy class parameterizing the set of controllers $\mathcal{K}_{\theta}$, where $\theta\in\mathbb{R}^{N}$ is a learnable parameter, which satisfies the following requirements by construction.
  1. (1)
    Robust stability certificates: All controllers in the policy class guarantee robust closed-loop stability with (1).
  2. (2)
    Differentiable parameterization: There exists a differentiable (at least piecewise) map from the learnable parameter $\theta\in\mathbb{R}^{N}$ to feedback policies $\mathcal{K}_{\theta}$.
  3. (3)
    Expressive policy class: The policy class is expressive and contains a large set of (preferably all) robustly stabilizing controllers for a given system (1).
Requirements (1) and (2) are hard requirements which are essential for learning-based control with robust stability guarantees in data-driven problems. Requirement (1) ensures the policies are always closed-loop stabilizing independently of the data or optimization algorithms with which they have been trained. Requirement (2) ensures that the policy class is easy to use and compatible with standard tools in machine learning and data-driven control which are based on unconstrained optimization. Requirement (3) is a soft requirement ensuring the policy class is sufficiently general to include high-performing controllers for a wide range of dynamical systems and objective functions. In a sense, our goal is to maximize requirement (3) subject to requirements (1) and (2).
With robust stability inherent to the policy class, we can learn stabilizing controllers $\mathcal{K}_{\theta}$ that seek to minimize arbitrary cost functions. A typical cost structure in deep RL is
$J=\mathop{\mathbb{E}}_{x_{0},\eta,d}\left[\sum_{t=0}^{T-1}g(x_{t},u_{t})+g_{f}(x_{T})\right]\quad\forall\,T\in\mathbb{N},$ (2)
where $g,g_{f}$ are any piecewise differentiable stage and terminal costs, respectively.
2.2 Definitions of Stability & Robustness
We now make precise the notions of stability and robustness that we will use. In contrast to linear time-invariant (LTI) systems, many different definitions of stability exist for nonlinear systems [33]. We use contraction [34] as a condition for internal stability, and incremental IQCs [35] for input-output stability and robustness.
Consider a nonlinear system $u\mapsto y$ of the form
$x_{t+1}=F(x_{t},u_{t}),\quad y_{t}=H(x_{t},u_{t}),$ (3)
with states $x_{t}\in\mathbb{R}^{n}$, inputs $u_{t}\in\mathbb{R}^{m}$, and outputs $y_{t}\in\mathbb{R}^{p}$. $F,H$ are assumed to be locally Lipschitz.
A contracting system is one in which all the state trajectories exponentially converge over time. That is, contracting systems forget their initial conditions exponentially [34, 49].
Definition 1 (Contraction)
The system (3) is said to be contracting if for any two initial states $x_{0}^{a},x_{0}^{b}\in\mathbb{R}^{n}$, given the same input sequence $u\in\ell^{m}$, the corresponding state trajectories $x^{a},x^{b}\in\ell^{n}$ satisfy
$|x^{a}_{t}-x^{b}_{t}|\leq\beta\alpha^{t}\absolutevalue{x_{0}^{a}-x_{0}^{b}}\quad\forall\,t\in\mathbb{N}$ (4)
with overshoot $\beta\in\mathbb{R}^{+}$ and contraction rate $\alpha\in[0,1)$.
We quantify the robustness of a system to variations in its inputs via incremental IQCs.
Definition 2 (Incremental IQC)
The system (3) is said to admit the incremental integral quadratic constraint (IQC) defined by $(Q,S,R)$ where $Q=Q^{\top}\in\mathbb{R}^{p\times p}$, $S\in\mathbb{R}^{m\times p}$, $R=R^{\top}\in\mathbb{R}^{m\times m}$, if for any two initial states $x^{a}_{0},x^{b}_{0}\in\mathbb{R}^{n}$ and input sequences $u^{a},u^{b}\in\ell^{m}$, the corresponding output sequences $y^{a},y^{b}\in\ell^{p}$ satisfy
$\sum_{t=0}^{T}\matrixquantity[\Delta y_{t}\\ \Delta u_{t}]^{\top}\matrixquantity[Q&S^{\top}\\ S&R]\matrixquantity[\Delta y_{t}\\ \Delta u_{t}]\geq-\kappa(x^{a}_{0},x^{b}_{0})\quad\forall T\in\mathbb{N}$ (5)
for some function $\kappa(x^{a}_{0},x^{b}_{0})\geq 0$ with $\kappa(x_{0},x_{0})=0\ \forall\,x_{0}\in\mathbb{R}^{n}$, and where $\Delta y_{t}=y^{b}_{t}-y^{a}_{t}$, $\Delta u_{t}=u^{b}_{t}-u^{a}_{t}$.
An important special case of (5) is when $Q=-\frac{1}{\gamma}$, $S=0$, $R=\gamma I$ for some $\gamma\in\mathbb{R}^{+}$, which defines a Lipschitz bound, a.k.a. incremental $\ell_{2}$-gain bound. Systems with small Lipschitz bounds are smooth in the sense that small changes to their inputs will not induce large variations in their outputs.
Definition 3 (Lipschitz system)
The system (3) is said to be Lipschitz if there exists a $\gamma\in\mathbb{R}^{+}$ such that
$\norm{\Delta y}_{T}\leq\gamma\norm{\Delta u}_{T}+\kappa(x^{a}_{0},x^{b}_{0})\quad\forall T\in\mathbb{N}$ (6)
with $\kappa(\cdot,\cdot),\Delta y,\Delta u$ as defined in Definition 2.
Many existing approaches to parameterizing stabilizing controllers are based on weaker notions of stability, such as finite $\ell_{2}$-gain stability. In the following definition, we assume $F(0,0)=0$ and $H(0,0)=0$ without loss of generality.
Definition 4 (Finite-gain stability)
The system (3) is said to be finite-gain stable if there exists a $\gamma\in\mathbb{R}^{+}$ such that for any initial state $x_{0}\in\mathbb{R}^{n}$, inputs $u\in\ell^{m}$, and corresponding outputs $y\in\ell^{p}$, we have
$\norm{y}_{T}\leq\gamma\norm{u}_{T}+\kappa(x_{0})\quad\forall T\in\mathbb{N}$ (7)
where $\kappa(x_{0})\geq 0$ and $\kappa(0)=0$.
3 Nonlinear Youla Parameterization
In this section, we construct the specific version of the Youla parameterization that we use. We assume that a stabilizing “base” controller has already been designed for the plant — or as a special case, where the plant itself is stable — and we wish to improve its closed-loop performance without compromising stability. Referring to Fig. 2, we learn a dynamical system $\mathcal{Q}$ called the Youla parameter which adds controls $\tilde{u}$ to the output of the base controller $\mathcal{K}_{b}$. The total control signal is then $u=\hat{u}+\tilde{u}$. We assume that $\mathcal{K}_{b}$ takes the form
$\mathcal{K}_{b}:\quad\left\{\begin{aligned} s_{t+1}&=f_{b}(s_{t},\eta_{t},u_{t},y_{t})\\ \hat{u}_{t}&=k(s_{t},\eta_{t},y_{t})\end{aligned}\right.$ (8)
where $s_{t}\in\mathbb{R}^{n_{s}}$ is some internal state. The first key ingredient in our parameterization is an observer $\hat{y}=\mathcal{O}(\eta,u,y)$ of the form
$\mathcal{O}:\quad\left\{\begin{aligned} \hat{x}_{t+1}&=f_{o}(\hat{x}_{t},\eta_{t},u_{t},y_{t}),\\ \hat{y}_{t}&=h(\hat{x}_{t})\end{aligned}\right.$ (9)
where $\hat{x}_{t}\in\mathbb{R}^{n_{x}}$ is the state estimate and $\hat{y}_{t}\in\mathbb{R}^{n_{y}}$ is the output estimate. The innovations sequence is
$\tilde{y}_{t}:=y_{t}-\hat{y}_{t}.$ (10)
Intuitively, the role of the observer is to decompose the measurements $y=\hat{y}+\tilde{y}$ into predictable measurements $\hat{y}$ and “surprises” $\tilde{y}$. The second key ingredient is the Youla parameter $\tilde{u}=\mathcal{Q}(\eta,\tilde{y})$, which acts on the innovations, “reacting to surprises” as well as known exogenous inputs $\eta$. $\mathcal{Q}$ is itself a nonlinear dynamical system of the form
$\mathcal{Q}:\quad\left\{\begin{aligned} q_{t+1}&=f_{q}(q_{t},\eta_{t},\tilde{y}_{t}),\\ \tilde{u}_{t}&=h_{q}(q_{t},\eta_{t},\tilde{y}_{t})\end{aligned}\right.$ (11)
with internal state $q_{t}\in\mathbb{R}^{n_{q}}$ and augmenting control $\tilde{u}_{t}\in\mathbb{R}^{n_{u}}$. The resulting augmented controller $u=\mathcal{K}_{\mathcal{Q}}(\eta,y)$ is
$\mathcal{K}_{\mathcal{Q}}:\quad\left\{\begin{aligned} s_{t+1}&=f_{b}(s_{t},\eta_{t},u_{t},y_{t})\\ \hat{x}_{t+1}&=f_{o}(\hat{x}_{t},\eta_{t},u_{t},y_{t})\\ q_{t+1}&=f_{q}(q_{t},\eta_{t},y_{t}-h(\hat{x}_{t}))\\ u_{t}&=k(s_{t},\eta_{t},y_{t})+h_{q}(q_{t},\eta_{t},y_{t}-h(\hat{x}_{t})).\end{aligned}\right.$ (12)
Figure 2: A version of the Youla-Kucera parameterization. The Youla parameter $\mathcal{Q}$ augments an observer-based feedback controller $\mathcal{K}_{b}$ to control a plant $\mathcal{G}$. The known exogenous inputs $\eta$ have been omitted to simplify the diagram.
Remark 1
If external inputs $\eta$ are not provided to the Youla parameter such that $\tilde{u}=\mathcal{Q}(\tilde{y})$, then the Youla parameter only reacts to surprises — i.e., disturbances, model error, or initial condition error. In this case, we can think of the controller (12) as consisting of a nominal controller $\mathcal{K}_{b}$ with good performance in response to known inputs, and a robustifying parameter $\mathcal{Q}$ whose role is to respond to error and uncertainty. This perspective is explored in the linear setting in [50].
4 Results for Simplified Settings
Our aim is to study the Youla controller $\mathcal{K}_{\mathcal{Q}}$ as a parameterization of policies for (a) nonlinear (b) partially-observed systems achieving (c) contracting and Lipschitz closed loops. In this section, we first consider simplified settings in which at most two of these three features are present at a time, namely: (b), (c) but not (a) (Sec. 4.1); (a), (c) but not (b) (Sec. 4.2); and (a), (b) but not (c) (Sec. 4.3), where the stability requirement is relaxed to finite-gain stability. In these settings, we find that a controller stabilizes the closed-loop system if and only if it can be written in the Youla form. Proofs of Theorems 1 and 2 are included as special cases of the results in Section 5. The proof of Theorem 3 is omitted due to space constraints, but similar results appear in [29, 30].
4.1 Restricting to LTI Systems
The following result from[40] is included here for completeness. Consider the simplified setting in which (1) is an LTI system
$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad y_{t}=Cx_{t}+v_{t}$ (13)
with $(A,B)$ stabilizable and $(A,C)$ detectable. Then, we can obtain a linear base controller $\hat{u}=\mathcal{K}_{b}(y)$ as
$\begin{split}\hat{x}_{t+1}&=A\hat{x}_{t}+Bu_{t}+L(y_{t}-C\hat{x}_{t})\\ \hat{u}_{t}&=-K\hat{x}_{t}\end{split}$ (14)
where the gain matrices $K,L$ are chosen such that $(A-BK)$ and $(A-LC)$ are stable. By taking the innovation $\tilde{y}=y-C\hat{x}$, the Youla controller (12) becomes
$u=-K\hat{x}+\mathcal{Q}(\eta,\tilde{y})$ (15)
where $\mathcal{Q}$ is the nonlinear system (11).
If $\mathcal{Q}$ is an LTI system, it is well-known that (15) parameterizes all linear stabilizing controllers for the plant (13) if and only if $\mathcal{Q}$ itself is stable [18, 16]. We extended this classic result in [40] to a nonlinear $\mathcal{Q}$ of the form (11) and nonlinear controllers $u=\mathcal{K}(\eta,y)$ with state-space realizations
$\phi_{t+1}=f_{\mathcal{K}}(\phi_{t},\eta_{t},y_{t}),\quad u_{t}=h_{\mathcal{K}}(\phi_{t},\eta_{t},y_{t}),$ (16)
states $\phi\in\mathbb{R}^{n_{k}}$, and $f_{\mathcal{K}},h_{\mathcal{K}}$ locally Lipschitz.
Theorem 1 (Prop. 1 from [40])
Consider the LTI system (13) with controller (14), (15) parameterized by $\mathcal{Q}$ (11).
  1. 1.
    For any contracting and Lipschitz $\mathcal{Q}$, the closed-loop system is contracting and the map $(\eta,d)\mapsto z$ is Lipschitz.
  2. 2.
    Any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a contracting and Lipschitz closed loop can be written in the form (14), (15) with contracting and Lipschitz $\mathcal{Q}$.
The Youla controller (12) therefore parameterizes all contracting and Lipschitz closed loops for a given LTI system.
4.2 Restricting to Perfect State Feedback
Next, we consider settings in which the plant (1) is nonlinear but we have perfect knowledge of the states such that
$x_{t+1}=f(x_{t},\eta_{t},u_{t})+w_{t},\quad y_{t}=x_{t}$ (17)
where $f(x_{-1},\eta_{-1},u_{-1}):=0$ and $w_{-1}:=x_{0}$. Suppose the base controller (8) takes the form $\hat{u}_{t}=k(\eta_{t},y_{t})$. Then the Youla parameterization (12) can be reduced to
$u=k(\eta,x)+\mathcal{Q}(\eta,\tilde{y})$ (18)
where the innovations at time $t$ are simply
$\tilde{y}_{t}=w_{t-1}=x_{t}-f(x_{t-1},\eta_{t-1},u_{t-1})$ (19)
since we have a perfect model of the plant and no measurement noise. A similar statement to Theorem 1 is then as follows.
Theorem 2
Suppose that $k$ is Lipschitz, the closed-loop system under the base controller $\hat{u}=k(\eta,x)$ is contracting, and the map $(\eta,w,\tilde{u})\mapsto z$ is Lipschitz. Consider the system (17) with controller (18) parameterized by $\mathcal{Q}$ (11).
  1. 1.
    For any contracting and Lipschitz $\mathcal{Q}$, the closed-loop system is contracting and its map $(\eta,d)\mapsto z$ is Lipschitz.
  2. 2.
    Any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a contracting and Lipschitz closed-loop can be written in the form (18) with contracting and Lipschitz $\mathcal{Q}$.
As with the LTI setting, the Youla controller (18) with innovations (19) parameterizes all contracting and Lipschitz closed-loops of (17). This formulation has strong connections to disturbance-feedback parameterizations used in MPC [21]. Similar results to Theorem 2 in terms of finite $\ell_{p}$-gain stability have also been derived in the context of systems-level synthesis (SLS) [51, 46].
4.3 Restricting to Finite-Gain Stability
Moving to the partially-observed nonlinear setting, similar results have been proven for the weaker notion of finite-gain stability (e.g., [29, 30]). We summarize these results for our specific problem setup below. Let us first make the following assumptions.
  1. A1
    Finite-gain stabilizing base controller: the closed-loop system $\mathcal{G}_{\mathcal{K}_{b}}:(\eta,d,\tilde{u})\mapsto z$ under the base controller, i.e., (1) in feedback with (8), is finite-gain stable.
  2. A2
    Contracting & Lipschitz observer: The observer (9) is contracting and the map $(\eta,u,y)\mapsto\hat{x}$ is Lipschitz.
  3. A3
    Observer correctness: When $d\equiv 0$ and $x_{0}=\hat{x}_{0}$, the observer exactly replicates the plant dynamics. That is, $f(x_{t},\eta_{t},u_{t})=f_{o}(x_{t},\eta_{t},u_{t},h(x_{t}))\ \forall t\in\mathbb{N}$.
  4. A4
    Smooth functions: The systems $\mathcal{G},\mathcal{K}_{b},\mathcal{O},\mathcal{Q},\mathcal{K}$ have piecewise-differentiable dynamics functions and Lipschitz continuous output maps.
Comparing to the partially-observed LTI setting, assumptions A1A3 are similar to assuming that we can design a stabilizing linear base controller (14). Importantly, A2 and A3 imply that the observer error will always exponentially decay in the absence of disturbances $d$, regardless of the control inputs $u$ [11, Prop. 2].
Under the above assumptions, the Youla parameterization achieves the following result (similar to e.g., [29, 30]).
Theorem 3
Suppose assumptions A1A4 hold. Consider the system (1) with controller (12) parameterized by $\mathcal{Q}$ (11).
  1. 1.
    For any finite-gain stable $\mathcal{Q}$, the closed-loop system is also finite-gain stable.
  2. 2.
    Any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a finite-gain stable closed loop can be written in the form (12) with finite-gain stable $\mathcal{Q}$.
This implies that the Youla parameterization represents all finite-gain stabilizing controllers for a given partially-observed nonlinear system.
5 Results for Partially-Observed Nonlinear Systems & Incremental Stability
In this section, we study the case of partially-observed nonlinear systems (1) and the incremental stability properties of contraction and Lipschitzness of the closed loop. We first show by counterexample that incremental stability can be lost in this context. We follow this by showing that a weaker-form of stability can be maintained, and discuss stronger conditions under which contracting and Lipschitz closed loops can still be achieved. All proofs are provided in Appendices .1.5.
We first introduce a stronger version of Assumption A1 which will be used for all subsequent results in this section
  1. A1
    Contracting & Lipschitz stabilizing base controller: the closed-loop system $\mathcal{G}_{\mathcal{K}_{b}}:(\eta,d,\tilde{u})\mapsto z$ under the base controller, i.e., (1) in feedback with (8), is contracting and Lipschitz.
5.1 Counterexample Exhibiting Loss of Contraction
One might expect a similar result to Theorems 13 to hold for partially-observed nonlinear systems in the incremental stability setting — i.e., if the base controller satisfies assumptions A1A4 and the Youla parameter is contracting and Lipschitz, then the Youla controller (12) parameterizes all contracting and Lipschitz closed loops $(\eta,d)\mapsto z$ for partially-observed nonlinear systems (1). However, Example 1 shows a system for which these assumptions are satisfied, but the closed-loop system is neither contracting nor Lipschitz.
Example 1
Consider the following scalar nonlinear system
$x_{t+1}=0.5\absolutevalue{x_{t}}+u_{t}+w_{t},\quad y_{t}=x_{t}$ (20)
with $w_{t}$ as the unknown disturbance. Assumption A1 holds with the base controller $\mathcal{K}_{b}=0$. We then choose an observer
$\hat{x}_{t+1}=0.5\absolutevalue{\hat{x}_{t}}+u_{t},\quad\hat{y}_{t}=\hat{x}_{t},$ (21)
which satisfies assumptions A2A4. Finally, we pass the innovations $\tilde{y}=y-\hat{y}$ to a memoryless Youla parameter
$u=\mathcal{Q}(\tilde{y}):=\min(2.5-5\tilde{y},0),$ (22)
which is obviously contracting and Lipschitz. As shown in Fig. 3 (green), the closed-loop state trajectories under the same disturbance $w\equiv 2$ converge to multiple trajectories, and hence the closed-loop system under the nonlinear Youla controller $\mathcal{K}_{\mathcal{Q}}$ is neither contracting nor Lipschitz.
Figure 3: Simulations of the closed-loop system in Example 1 show plant state trajectories from different initial states (solid and dashed). They converge when $w=0$ but with $w=2$ there are distinct trajectories that do not converge, so the system is neither contracting nor Lipschitz.
To understand the behavior of the system (20)–(22), we first derive the dynamics of the observer error $\tilde{x}=x-\hat{x}$ as follows,
$\tilde{x}_{t+1}=0.5|\tilde{x}_{t}+\hat{x}_{t}|-0.5|\hat{x}_{t}|+w_{t},\quad\tilde{y}_{t}=\tilde{x}_{t},$ (23)
which implies that the innovations $\tilde{y}$ implicitly depend on the control input $u$ (the output of $\mathcal{Q}$) through $\hat{x}$. This introduces a feedback loop between the error dynamics (23) and the Youla parameter $\mathcal{Q}$ (22), leading to non-contracting closed-loop behavior when the gain of $\mathcal{Q}$ is sufficiently large.
On the other hand, we can establish that increasing the gain of $\mathcal{Q}$ does not lead to unbounded closed-loop trajectories with bounded disturbances. The reason is as follows. Both the plant (20) and its observer (21) are copies of the following contracting and Lipschitz system
$\xi_{t+1}=0.5\absolutevalue{\xi_{t}}+\mu_{t},\quad\nu_{t}=\xi_{t}$ (24)
with $(\xi^{a},\mu^{a},\nu^{a})=(x,u+w,y)$ and $(\xi^{b},\mu^{b},\nu^{b})=(\hat{x},u,\hat{y})$ for the plant and observer, respectively. From (6) we have
$\|\tilde{y}\|_{T}=\|y-\hat{y}\|_{T}\leq\gamma\|w\|_{T}+\kappa(x_{0},\hat{x}_{0})$ (25)
with $\gamma>0$ and $\kappa(a,b)\geq 0$. In other words, the innovations $\tilde{y}$ are bounded by the external disturbance $w$, despite there being feedback through $u$, which might be large due to some high-gain $\mathcal{Q}$. Moreover, since the Youla parameter $\mathcal{Q}$ and the system under the base controller are both contracting and Lipschitz, we can conclude that the $\ell_{2}$-norms of $u,x,y$ are also bounded.
5.2 d-Tube Contracting and Lipschitz Responses
The counterexample above shows that closed-loop contraction and Lipschitzness cannot be guaranteed for partially-observed nonlinear systems (1). However, we observed that signals did not exhibit unbounded growth — they satisfied a finite $\ell^{2}$ gain bound from the unknown disturbance $w$. The main result of this section establishes this as a general property in the partially-observed nonlinear setting. In the following, $\bar{x}_{t}:=[x_{t};s_{t};\hat{x}_{t};q_{t}]$ is the state of the closed-loop system under the Youla controller (12).
Theorem 4
Suppose Assumptions A1A4 hold. Then for any contracting and Lipschitz Youla parameter $\mathcal{Q}$ (11), the closed-loop system consisting of (1) in feedback with (12) satisfies the following properties:
  1. 1.
    With zero disturbances (i.e., $d\equiv 0$, $\tilde{x}_{0}=0$), the closed-loop system is contracting and $\eta\mapsto z$ is Lipschitz.
  2. 2.
    Let $(\bar{x}^{\star},\eta^{\star},z^{\star})\in\ell$ be any zero-disturbance trajectory of the closed-loop system. Then with nonzero disturbances (i.e., $\forall d\in\ell$, $x_{0}\neq\hat{x}_{0}$), the closed-loop system satisfies
    $\norm{z-z^{\star}}_{T}\leq\gamma(\norm{\eta-\eta^{\star}}_{T}+\norm{d}_{T})+\kappa(\bar{x}_{0},\bar{x}^{\star}_{0})$ (26)
    for all $T\in\mathbb{N}$, $\bar{x}_{0}\in\mathbb{R}^{n_{\bar{x}}}$, $\eta\in\ell$, where $\gamma\in\mathbb{R}^{+}$, $\kappa(\bar{x}_{0},\bar{x}^{\star}_{0})\geq 0$, $\kappa(\bar{x}^{\star}_{0},\bar{x}^{\star}_{0})=0$, and $n_{\bar{x}}=2n_{x}+n_{q}+n_{s}$. We call this property d-tube contracting and Lipschitz.
The interpretation of Theorem 4 is that in the disturbance-free case, the closed-loop system exhibits smooth responses to known external inputs $\eta$ (e.g., disturbance previews, reference trajectories, feedforward controls) as they can be seen by both the plant and the observer. However, with unknown disturbances present, the system contracts to a tube around the disturbance-free trajectory with a radius proportional to $\|d\|$. This is a stronger property than the finite-gain stability of Theorem 3, but weaker than the closed-loop contraction and Lipschitzness of Theorems 1 and 2.
Remark 2
The proof of Theorem 4 (Appendix .1) does not actually rely on the base controller achieving contracting and Lipschitz closed loops w.r.t. $d$. A1 can be replaced with the weaker assumption that the base-controlled closed-loop system $\mathcal{G}_{\mathcal{K}_{b}}:(\eta,d,\tilde{u})\mapsto z$ is d-tube contracting and Lipschitz.
5.3 Response to Observer Initial Condition Error
Theorem 4 indicates that the Youla parameterization guarantees contracting and Lipschitz closed-loops in the disturbance-free setting with zero initial observer error, but not in the presence of disturbances and initial observer error. An interesting special case is when we have nonzero initial observer error, but no other unknown perturbations.
Consider Example 1 again, now with $w\equiv 0$. Since both $(x,u,y)$ and $(\hat{x},u,\hat{y})$ are solutions of the contracting system (24) with the same inputs $u$, from (4) we have $|\tilde{x}_{t}|=|x_{t}-\hat{x}_{t}|\leq\beta\alpha^{t}|x_{0}-\hat{x}_{0}|$ for some $\beta>0$ and $\alpha\in[0,1)$. This means that the feedback from $\mathcal{Q}$ to $\tilde{y}$ is constrained within an envelope that exponentially converges to zero. Since the Youla parameter $\mathcal{Q}$ and the system under the base controller $\mathcal{K}_{b}$ are both contracting, we conclude that both $x,u$ will exponentially converge to some fixed values, which we see empirically in Fig. 3 (yellow). The following theorem formalizes the closed-loop stability result when the unknown disturbances are zero.
Theorem 5
Suppose assumptions A1A4 hold, $x_{0}\neq\hat{x}_{0}$ but $d\equiv 0$, and $f,f_{b},f_{q}$ are Lipschitz continuous w.r.t. their respective system inputs. Consider any two trajectories $z^{a},z^{b}\in\ell$ of the closed-loop system (1) in feedback with (12) with initial states $\bar{x}^{a}_{0},\bar{x}^{b}_{0}\in\mathbb{R}^{n_{\bar{x}}}$ and inputs $\eta^{a},\eta^{b}\in\ell$. Then for any contracting and Lipschitz Youla parameter $\mathcal{Q}$ (11):
  1. 1.
    When $\eta^{a}\equiv\eta^{b}$, the closed-loop states satisfy
[…]
    for some $A\in\mathbb{R}^{+}$ and $\rho\in[0,1)$, where $\Delta x:=x^{a}-x^{b}$ and similar for all other variables, and $\xi_{t}:=[x_{t};s_{t}]$.
  2. 2.
    The closed-loop trajectories satisfy
    $\norm{\Delta z}_{T}\leq\bar{\gamma}\norm{\Delta\eta}_{T}+\kappa(\bar{x}^{a}_{0},\bar{x}^{b}_{0})+\frac{\tilde{\gamma}\beta\left(\absolutevalue{\tilde{x}^{a}_{0}}+\absolutevalue{\tilde{x}^{b}_{0}}\right)}{\sqrt{1-\alpha^{2}}}$ (28)
    for all $T\in\mathbb{N}$, where $\bar{\gamma},\tilde{\gamma}\in\mathbb{R}^{+}$, $\kappa(\bar{x}^{a}_{0},\bar{x}^{b}_{0})\geq 0$ with $\kappa(\bar{x}_{0},\bar{x}_{0})=0\ \forall\,\bar{x}_{0}\in\mathbb{R}^{n_{\bar{x}}}$, and where $\beta\in\mathbb{R}^{+},\alpha\in[0,1)$ are the overshoot and contraction rate of the observer (9).
Remark 3
The effect of nonzero initial observer error is that the closed-loop system is not contracting and Lipschitz in the sense of Definitions 1 and 3 (respectively). Instead, (27) and (28) have additional terms depending on $|\tilde{x}_{0}^{a}|,|\tilde{x}_{0}^{b}|$, which are not necessarily zero even if $\Delta\tilde{x}_{0}=0$. These terms are due to the transient feedback interaction between $\tilde{u}$ and $\tilde{y}$ before the observer error $\tilde{x}$ converges to 0. We therefore called this behavior contracting and Lipschitz with transients in [43], and it is consistent with recent results in [45, Thm. 4.3].
5.4 Decoupled Innovations & Youla Augmentation
The source of the loss of contraction in the counterexample (Sec. 5.1) was the indirect interaction between the augmentation input $\tilde{u}$ and the innovations $\tilde{y}$. If, however, $\tilde{y}$ is completely decoupled from $\tilde{u}$, then adding $\mathcal{Q}$ does not introduce a feedback loop in the closed-loop system, and we retain the closed-loop contracting and Lipschitz properties of the original base-controlled system. We have already seen this property in three special cases — LTI systems (Thm. 1.1), fully-observed nonlinear systems (Thm. 2.1), and partially-observed nonlinear systems with zero uncertainty (Thm. 4.1).
Proposition 1
In each of the closed-loop systems from Theorems 1, 2, and 4.1, the innovations $\tilde{y}$ is the output of a contracting and Lipschitz system $\mathcal{T}:(\eta,d)\mapsto\tilde{y}$ given by
$\psi_{t+1}$ $=\tilde{F}(\psi_{t},\eta_{t},d_{t})$ (29)
$\tilde{y}_{t}$ $=\tilde{H}(\psi_{t},\eta_{t},d_{t})$
with internal state $\psi_{t}\in\mathbb{R}^{n_{\psi}}$ and $\tilde{F},\tilde{G}$ locally Lipschitz.
The decoupling of $\tilde{y}$ and $\tilde{u}$ in this way is a sufficient condition for the Youla parameterization to guarantee contracting and Lipschitz closed-loop responses. We therefore unify our three special cases with the following result.
Theorem 6
Suppose the base-controlled system $\mathcal{G}_{\mathcal{K}_{b}}:(\eta,d,\tilde{u})\mapsto z$ satisfies assumption A1 and the innovations $\tilde{y}$ are generated by a contracting and Lipschitz system $\mathcal{T}$ (29). Then for any contracting and Lipschitz Youla parameter $\mathcal{Q}$ (11), the closed-loop system $z=\mathcal{G}_{\mathcal{K}_{b}}(\eta,d,\mathcal{Q}(\eta,\mathcal{T}(\eta,d)))$ mapping $(\eta,d)\mapsto z$ is also contracting and Lipschitz.
5.5 Achieving Closed-Loop Contracting & Lipschitz Responses to Disturbances and Uncertainty
Figure 4: Youla parameterization via loop shaping, where known inputs $\eta$ have been omitted for simplicity.
When the innovations are coupled with the Youla augmentation (e.g., due to disturbances or model parameter uncertainty), we can still achieve contracting and Lipschitz closed loops by imposing further restrictions on the Youla parameter. Our approach is based on IQC theory and loop shaping. We first group together the nonlinear plant $\mathcal{G}$, the base controller $\mathcal{K}_{b}$, and the observer $\mathcal{O}$ as the system $\mathcal{G}_{\Delta}:(\eta,d,\tilde{u})\mapsto(z,\tilde{y}$) shown in Fig. 4. We then take $\mathcal{Q}=\mathcal{W}_{1}\circ\bar{\mathcal{Q}}\circ\mathcal{W}_{2}$ where $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ are fixed, stable linear filters, and $\bar{\mathcal{Q}}$ is a learnable contracting and Lipschitz system. $\mathcal{Q}$ is therefore also contracting and Lipschitz. The following proposition gives a sufficient condition for achieving a stable closed loop.
Proposition 2
Suppose the system $\mathcal{G}_{\Delta}$ is contracting and there exist stable linear filters $\mathcal{W}_{1}:\bar{u}\mapsto\tilde{u}$ and $\mathcal{W}_{2}:\tilde{y}\mapsto\bar{y}$ with $\bar{u}\in\ell^{n_{u}}$ and $\bar{y}\in\ell^{n_{y}}$ such that the weighted system $\bar{\mathcal{G}}_{\Delta}:(\eta,d,\bar{u})\mapsto(z,\bar{y})$ admits the incremental IQC defined by $(Q,S,R)$ with $Q\prec 0$ and $R\succ 0$ partitioned as
$Q=\begin{bmatrix}Q_{zz}&Q_{z\bar{y}}\\ Q_{z\tilde{y}}^{\top}&Q_{\bar{y}\bar{y}}\end{bmatrix},S=\begin{bmatrix}S_{dz}&S_{d\bar{y}}\\ S_{\bar{u}z}&S_{\bar{u}\bar{y}}\end{bmatrix},R=\begin{bmatrix}R_{dd}&R_{d\bar{u}}\\ R_{d\bar{u}}^{\top}&R_{\bar{u}\bar{u}}\end{bmatrix}.$
Take $\mathcal{Q}=\mathcal{W}_{1}\circ\bar{\mathcal{Q}}\circ\mathcal{W}_{2}$, and suppose further that the system $\bar{\mathcal{Q}}:\bar{y}\mapsto\bar{u}$ is contracting and satisfies the incremental IQC defined by $(\bar{Q},\bar{S},\bar{R})$. Then the closed loop of $\mathcal{G}_{\Delta}$ and $\mathcal{Q}$ is contracting and Lipschitz if
$\begin{bmatrix}Q_{\bar{y}\bar{y}}+\bar{R}&S_{\bar{u}\bar{y}}^{\top}+\bar{S}&Q_{z\bar{y}}\\ S_{\bar{u}\bar{y}}+\bar{S}^{\top}&R_{\bar{u}\bar{u}}+\bar{Q}&S_{\bar{u}z}\\ Q_{z\bar{y}}^{\top}&S_{\bar{u}z}^{\top}&Q_{zz}\end{bmatrix}\prec 0.$ (30)
Remark 4
Using a pre-compensator $\mathcal{W}_{1}$ and/or a post-compensator $\mathcal{W}_{2}$, the system $\tilde{\mathcal{G}}_{\Delta}$ has a desired open-loop shape, which can be understood as a dynamic version of input normalization in deep RL [52]. This allows for learning high-performance controllers via simple IQCs for $\bar{\mathcal{G}}_{\Delta}$ and $\bar{\mathcal{Q}}$, e.g.,
$\begin{split}Q_{\tilde{y}\tilde{y}}=-1/\nu I,\ &S_{\tilde{u}\tilde{y}}=0,\ R_{\tilde{u}\tilde{u}}=\nu I,\ Q_{z\bar{y}}=0,\ S_{\bar{u}z}=0,\\ &\bar{Q}=-1/\mu I,\ \bar{S}=0,\ \bar{R}=\mu I.\end{split}$
In this case, condition (30) is reduced to the requirement for the incremental small-gain theorem, i.e., $\mu\nu<1$.
6 Converse Results
In Section 4 we saw that, in certain simplified settings, the proposed Youla parameterization includes all and only controllers resulting in contracting and Lipschitz closed-loops. In Section 5 we saw by counterexample that this is not true in the partially-observed nonlinear case, but that the weaker property of d-tube contraction and Lipschitzness holds. This raises the question: what class of controllers are covered by the Youla parameterization? We provide an answer in this section, and include proofs in Appendix .6.
We first introduce the following system which connects the observer (9) and base controller dynamics (8) in parallel, forming a map $(\eta,u,\tilde{y})\mapsto(y,\hat{z})$:
$\mathcal{O}:\hat{x}_{t+1}$ $=f_{o}(\hat{x}_{t},\eta_{t},u_{t},\hat{y}_{t}+\tilde{y}_{t}),\quad\hat{y}=h(\hat{x}_{t})$ (31)
$\mathcal{K}_{b}:s_{t+1}$ $=f_{b}(s_{t},\eta_{t},u_{t},\hat{y}_{t}+\tilde{y}_{t}),\quad\hat{u}=k(s_{t},\eta_{t},\hat{y}_{t}+\tilde{y}_{t})$
$y_{t}$ $=h(\hat{x}_{t})+\tilde{y}_{t}$
$\hat{z}_{t}$ $=[\hat{x}_{t};s_{t};u_{t}]$
We then introduce an analogous assumption to A1 for (31).
  1. A1′′
    Observer-stabilizing base controller: the closed-loop system $\tilde{\mathcal{G}}_{\mathcal{K}_{b}}:(\eta,\tilde{y},\tilde{u})\mapsto\hat{z}$ under the base controller, i.e., (31) in feedback with (8), is contracting and Lipschitz.
Theorem 7
Suppose assumption A4 holds, and consider the system (31) with controller (12) parameterized by $\mathcal{Q}$ (11).
  1. 1.
    Suppose assumption A1′′ also holds. Then for any contracting and Lipschitz $\mathcal{Q}$, the closed-loop system with (31) is contracting and the map $(\eta,\tilde{y})\mapsto\hat{z}$ is Lipschitz.
  2. 2.
    Any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a contracting and Lipschitz closed loop with (31) can be written as (12) with a contracting and Lipschitz $\mathcal{Q}$.
The interpretation is that for a given base controller and observer, the Youla controller (12) with contracting and Lipschitz $\mathcal{Q}$ parameterizes all and only controllers $\mathcal{K}$ stabilizing the system (31). The structure of this interconnection is shown in Fig. 5. It can be seen that there is a feedback interconnection of the controller $\mathcal{K}$ and the observer $\mathcal{O}$ (red), which is connected in series/parallel with the base controller $\mathcal{K}_{b}$ (green).
Figure 5: Block diagram showing how a Youla parameter $\tilde{u}=\mathcal{Q}(\tilde{y})$ can be constructed so that any controller $u=\mathcal{K}(y)$ can be realized in the Youla form (12). Known exogenous inputs $\eta$ have been omitted to simplify the diagram.
The proof is based on the idea that for a given base controller and observer, any controller $\mathcal{K}$ can be realized in the Youla form (12) with $\mathcal{Q}$ as the mapping $(\eta,\tilde{y})\mapsto\tilde{u}$ shown in Fig. 5. With this construction, the mapping $y\mapsto u$ of the controller $\mathcal{K}$ remains unchanged. Stability of the resulting map $(\eta,\tilde{y})\mapsto\hat{z}$ is essentially equivalent to stability of each of the red and green components, since $\hat{z}$ includes the states of both, and contraction and Lipschitzness are preserved through series/parallel system connections. We take a deeper look into the meaning of stabilizing the red and green systems below.
1) Stabilizing the red system corresponds to the controller $\mathcal{K}$ stabilizing the observer. This is related to stabilizing the plant (1), but not the same when disturbances are present. When $\tilde{y}=0$, observer correctness (A3) implies that $f_{o}(x_{t},\eta_{t},u_{t},h(x_{t}))=f(x_{t},\eta_{t},u_{t})$, which is exactly the disturbance-free plant dynamics. The controller $\mathcal{K}$ stabilizes the observer against the particular disturbance structure
$\hat{x}_{t+1}=f_{o}(\hat{x}_{t},\eta_{t},u_{t},h(x_{t})+w_{t})$
entering via the measurement channel, rather than simply an additive disturbance as in (1).
2) Stability of the green system is just stability of the base controller. At first, this may look like a requirement of strong stabilization, which is known to be restrictive. However, this is not the case since $\mathcal{K}_{b}$ takes the true control signal $u$ as an input. For example, in the LTI case from Sec. 4.1, this is the requirement that $(A-LC)$ is stable, which will generally be the case, whereas strong stabilization requires that $(A-LC-BK)$ be stable, which is restrictive. The difference is that we use the total control signal $u$ in place of $\hat{u}=-K\hat{x}$.
Remark 5
Stabilizing the red system in Fig. 5 can be interpreted as defining a stable image representation of the controller $\mathcal{K}$ (i.e., a right-coprime factor), since it represents a mapping from $\tilde{y}\mapsto(\mathcal{K}(y),y)$. Similarly, one can think of the original observer (9) together with the innovations map (10) as a stable kernel representation (i.e., a left-coprime factor) mapping $(u,y)\mapsto\tilde{y}$ (where we have omitted $\eta$ for illustrative purposes). The perspective of constructing the Youla parameterization via stable image and kernel representations is explored in detail in [29, 45].
In certain special cases, the requirement that a controller $\mathcal{K}$ stabilizes the system (31) follows naturally from $\mathcal{K}$ stabilizing the plant (1). These special cases arise when we make stronger assumptions on the structure of the plant, base controller, and observer. In the following corollaries, we show that the Youla parameterization contains all contracting and Lipschitz closed-loops for two particular system structures. Note that Theorem 1.2 for LTI systems follows directly from Corollary 2.
Corollary 1
Theorem 2.2 follows directly from Theorem 7 with measurements $y_{t}=x_{t}$, observer $\hat{x}_{t+1}=f(y_{t},\eta_{t},u_{t})$, and a static base controller $\hat{u}_{t}=k(\eta_{t},y_{t})$.
Corollary 2
Suppose assumption A4 holds and the base controller (8) and observer (9) are in the certainty-equivalence form
$\hat{x}_{t+1}$ $=f(\hat{x}_{t},\eta_{t},u_{t})+L(\eta_{t},y_{t}-h(\hat{x}_{t}))$ (32)
$\hat{u}_{t}$ $=k(\hat{x}_{t},\eta_{t},y_{t}),$
where $L:\mathbb{R}^{n_{\eta}}\times\mathbb{R}^{n_{y}}\rightarrow\mathbb{R}^{n_{x}}$ is Lipschitz. Then any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a contracting and Lipschitz closed loop $(\eta,d)\mapsto z$ with (1) can be written in the form (12) with a contracting and Lipschitz $\mathcal{Q}$ (11).
Remark 6
To understand why Corollary 2 cannot be extended to a necessary and sufficient condition parameterizing all and only stabilizing controllers $\mathcal{K}$, we examine the difference between stabilizing the system (31) and stabilizing the plant (1). Consider Example 1 again. In this case, the observer (21) is just an open-loop simulation of the disturbance-free plant (20). It does not depend on $y$ (or $\tilde{y}$) at all, hence adding a disturbance to the “measurement channel” of the observer has no effect (i.e., the Lipschitz constant of $f_{o}$ w.r.t. $y$ is zero). Therefore, a controller $\mathcal{K}$ stabilizing this observer only stabilizes the plant in a disturbance-free setting, which implies nothing about its ability to stabilize the plant in the presence of additive disturbances $w$.
Fig. 6 illustrates the connection between Theorems 4, 6, and 7 for stabilizing controllers satisfying the conditions in Theorem 7. The Youla parameterization (12) contains all policies (16) that achieve contracting and Lipschitz closed loops with the plant (set C). When the innovations and control inputs are coupled due to additive disturbances, the closed-loop system does not necessarily achieve these same strong stability properties. Instead, it guarantees d-tube contraction and Lipschitzness, which is a slightly weaker stability result (set A). In the special case where the innovations are decoupled from the control signal as per (29), all three sets coincide (e.g., LTI systems, full state-feedback).
Figure 6: Illustration of stabilizing policy sets for controllers satisfying the conditions in Theorem 7: A) controllers achieving d-tube contracting and Lipschitz closed-loops, B) the proposed nonlinear Youla parameterization, and C) controllers achieving contracting and Lipschitz closed loops. The three sets coincide if the innovations also satisfy (29).
7 Deep RL with Youla-REN
Having established the stability properties of the nonlinear Youla parameterization (12), we now investigate its use in learning-based control tasks. In this section, we first introduce our proposed Youla-REN policy class (Sec. 7.1), where we use a REN [11] as the contracting and Lipschitz Youla parameter $\mathcal{Q}$. We then illustrate the advantages of its built-in stability certificates on three deep RL tasks with: (i) general reward functions without stability consideration; (ii) short training horizons; and (iii) systems with model uncertainty (Sec. 7.27.4, respectively). All experiments [^1] https://github.com/nic-barbara/ReactToSurprises were performed in Julia using the RobustNeuralNetworks.jl package [53] and the analytic policy gradients RL algorithm [14].
7.1 The Youla-REN Policy Class
We parameterize the Youla parameter $\mathcal{Q}$ via recurrent equilibrium networks (RENs), which are nonlinear systems
[…]
with learnable parameters $W,b$, where $q_{t}\in\mathbb{R}^{n_{q}}$ is the internal state, and $\bar{v}_{t},\bar{w}_{t}\in\mathbb{R}^{n_{\bar{v}}}$ are the input and output of the neural layer, respectively. Here $\sigma$ is a scalar activation with its slope restricted in $[0,1]$ (e.g., relu or tanh). Together with (12), the resulting policy class is called Youla-REN. The motivations for choosing a REN as the Youla parameter are as follows.
  1. 1.
    RENs admit built-in behavioral guarantees such as contraction, Lipschitzness, or other properties described by incremental IQCs. Thus, they are compatible with the analysis framework for our policy parameterization. In the experiments below, we use $\gamma$REN to denote a REN with a prescribed Lipschitz bound of $\gamma\in\mathbb{R}^{+}$.
  2. 2.
    RENs are highly flexible and include many existing models as special cases, including multi-layer perceptrons, recurrent neural networks, and stable linear dynamical systems. It was proven in [40, Prop. 2] that contracting RENs are also universal approximators for contracting and Lipschitz dynamical systems.
  3. 3.
    RENs are compatible with standard machine-learning tools as they permit a direct (smooth, unconstrained) parameterization $\theta\mapsto(W,b)$ such that for any $\theta\in\mathbb{R}^{N}$, the resulting REN $\mathcal{Q}_{\theta}$ (33) is contracting and Lipschitz. This enables training of large-scale models via automatic differentiation and variants of stochastic gradient descent.
7.2 Nonlinear System with Economic Cost
In our first example, we learn stabilizing policies for partially-observed nonlinear system while optimizing an economic cost that does not naturally encourage stability. Consider the following academic example, adapted from [54]:
$\dot{x}_{1}$ $=-x_{1}+x_{3}+w_{1}$ (34)
$\dot{x}_{2}$ $=x_{1}^{2}-x_{2}-2x_{1}x_{3}+x_{3}+w_{2}$
$\dot{x}_{3}$ $=-x_{2}+u+w_{3}$
$y$ $=[x_{2};x_{3}]+v.$
We chose a stabilizing base controller $u=y_{1}-ky_{2}$ with gain $k=1.5$, and an observer of the form
$\dot{\hat{x}}_{1}$ $=-\hat{x}_{1}+\hat{x}_{3}$ (35)
$\dot{\hat{x}}_{2}$ $=\hat{x}_{1}^{2}-\hat{x}_{2}-2\hat{x}_{1}\hat{x}_{3}+\hat{x}_{3}$
$\dot{\hat{x}}_{3}$ $=-\hat{x}_{2}+u+(y_{2}-\hat{x}_{3})-(y_{1}-\hat{x}_{2}).$
It can be shown that Assumptions A1A4 hold for the this base controller and observer via [55]. We set $w,v$ as zero-mean Gaussian noise with variances $\sigma_{w_{i}}=10^{-2},\sigma_{v}=10^{-3}$ respectively, and discretized the system using the forward-Euler approximation and a time-step of $\Delta t=0.05$ s.
(a) Test cost during training.
(b) Trajectory rollouts of $u$ and $x_{2}$ for a trained Youla-REN.
Figure 7: Youla-REN learns to improve economic performance while maintaining closed-loop stability. Bands in (a) show the range across 10 random seeds. Dotted lines in (b) indicate the soft constraints ±$\pm u_{\mathrm{max}}$ imposed on the controls.
We chose an economic stage cost
$g(x_{t},u_{t}):=|u_{t}|+500\max(|u_{t}|-u_{\mathrm{max}},0),$ (36)
which is the $\ell_{1}$-norm of the control inputs plus a strong penalty for exceeding the soft threshold $u_{\mathrm{max}}$. Note that if stability is ignored, then the optimal policy is simply $u\equiv 0$, which results in an unstable closed loop.
Fig. 7 shows that with the proposed Youla-REN, we are able to learn a stabilizing controller which achieves significant performance improvements compared to the base controller while maintaining closed-loop stability. The learned Youla parameter $\mathcal{Q}$ adds corrections to the base controller signal to keep the total $u$ within the soft bounds ±$\pm u_{\mathrm{max}}$ most of the time, allowing only small deviations outside the bounds for short intervals, and encouraging it close to zero within these bounds. Moreover, the trajectories of both $u$ and $x_{2}$ indicate that the Youla-REN preserves the closed-loop stability of the base controller even though such behavior was not explicitly encoded in the cost function. We therefore avoid the laborious process of reward-shaping to encourage stability in this task, which is commonly required for deep RL policies to be both stabilizing and performant (e.g., [2]).
7.3 Long-Term Stability with Short-Term Training
Our second example illustrates that our proposed Youla-REN policies can achieve good long-term test performance even when they are trained over short time horizons. We consider the following well-known linear quadratic Gaussian (LQG) example from [56],
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}$ $=\matrixquantity[1\hfil&1\\ 0&1]\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}+\matrixquantity[0\\ 1]u+\matrixquantity[1\\ 1]w$ (37)
$y$ $=x_{1}+v,$
where $w,v\in\mathbb{R}$ are Gaussian noises with variances $\sigma_{w}>0$ and $\sigma_{v}=1$, respectively. The control objective is
[…]
with $q>0$. The optimal LQG controller has arbitrarily small gain and phase margins as $\sigma_{w},q$ increase [56]. This makes the problem challenging in a deep RL setting, since optimal and unstable policies may appear very close in parameter space.
We discretized the system with the forward-Euler method ($\Delta t=0.01$ s) and chose a training horizon of 100 s. For the base controller, we chose a de-tuned LQG controller designed to minimize (38) with $q=10^{3}$ and $\sigma_{w}=10$, while for training and test evaluations we used $q,\sigma_{w}=10^{3}$.
(a) Mean training cost based on short-horizon rollouts.
(b) Mean test cost for long-term policy rollouts.
Figure 8: Youla-REN achieves long-term stability even with short-horizon training on the classic LQG problem from [56]. Bands in (a) show the range across 10 random seeds.
For comparison, we also trained Residual-REN policies
$u=\mathcal{K}_{b}(y)+\mathcal{Q}(y),$ (39)
where $\mathcal{K}_{b}$ is the stabilizing base controller and the learnable component $\mathcal{Q}$ is parameterized by a REN. As shown in Fig. 1, the main difference from Youla-REN is that the system $\mathcal{Q}$ in a Residual-REN acts directly on measurements $y$ rather than the innovations $\tilde{y}$. A direct consequence is that Youla-REN guarantees closed-loop stability (provided that $\mathcal{Q}$ is contracting and Lipschitz) while Residual-REN does not.
The training curves in Fig. 8(a) show that the Youla-REN policies consistently outperform the Residual-RENs and smoothly converge to a near-optimal final cost. In contrast, the Residual-RENs never approach the optimal LQG cost and yield large variance during training. We found that the Residual-RENs were highly sensitive to the learning rate, with slightly smaller values resulting in almost no improvement over the base controller, while larger values lead to even greater fluctuations and a worse final cost. Since the true optimal policy is near the boundary of the stabilizing controller set, it is common to encounter unstable Residual-REN policies during training (e.g., at epoch 687 in the top panel of Fig. 8(b)). This is not possible with the Youla-REN due to its built-in stability guarantees.
We also compare the long-term test performance of each policy architecture in Fig. 8(b). Testing Residual-RENs at epochs 1500 and 4000, the policies appears to stabilize the closed-loop system over the training horizon (i.e., the cost is less than the base cost). However, rolling out these policies over longer horizons reveals unstable responses. The Youla-REN demonstrates consistently strong performance over long-term horizons at all stages during and after training.
7.4 Robust Stability with Model Uncertainty
Our final example demonstrates that Youla-REN policies can achieve good performance while maintaining closed-loop stability in the presence of model parameter uncertainty. Consider the linearized cart-pole system
$\dot{x}$ $=\matrixquantity[0\hfil&1&0&0\\ 0&0&-\frac{m_{p}g}{m_{c}}0\\ 0&0&0&1\\ 0&0&\frac{(m_{c}+m_{p})g}{m_{c}l}0]x+\matrixquantity[0\\ \frac{1}{m_{c}}\\ 0\\ -\frac{1}{m_{c}}]u+w$ (40)
$y$ $=\matrixquantity[1\hfil&0&0&0]x+v$
where $w,v$ are zero-mean Gaussian noise with variances $\sigma_{w}=10^{-3}$, $\sigma_{v}=10^{-4}$, respectively. We discretized (40) with the forward-Euler method ($\Delta t=0.02$ s) and chose $g=9.81$, $l=10$, $m_{c}=1$ and $m_{p}\in[0.14,0.35]$. The control objective was to minimize an infinite-horizon quadratic cost with weights $Q=\texttt{diag}(1,0,1,0)$ and $R=I$. We chose a discrete-time, time-invariant LQG controller as the base controller, designed at a nominal mass $m_{p}^{\star}=0.2$. We computed its control gain with the same $Q,R$ as the cost function, but the steady-state Kalman filter was designed with covariances $\Sigma_{w}=\texttt{diag}(1,10^{3},1,10^{3})\times\Delta t$ and $\Sigma_{v}=10^{-3}I/\Delta t$ for the process and measurement noises, respectively, to ensure closed-loop stability across the whole uncertain $m_{p}$ range.
(a) $\tilde{u}\mapsto\tilde{y}$
(b) $\tilde{u}\mapsto y$
Figure 9: Bode magnitude plots for system (40) under the base LQG controller with different pole masses $m_{p}$, where the color gradient darkens as $m_{p}$ increases. The yellow lines show the inverse of the (discretized) weighting filters (41).
(a) Training curves.
(b) Cost rollouts.
(c) Sensitivity to model uncertainty.
Figure 10: Mean test cost for the cart-pole problem with uncertain pole mass: (a) short-horizon cost during training (2 times training horizon); (b) cost for long-term rollouts of particular policies; and (c) long-horizon cost as a function of the uncertain pole mass (8 times training horizon). LQG refers to the optimal time-varying implementation while LTI LQG refers to the combination of steady-state Kalman filter and LQR gain.
We parameterized robustly-stabilizing policies for this system with Youla-$\gamma$REN, a Lipschitz-bounded variant of Youla-REN. We chose the Lipschitz bound $\gamma$ using a small-gain approach with loop shaping to enable both closed-loop stability and strong performance (Prop. 2). For frequency weightings, we chose the output filter to be $\mathcal{W}_{1}(s)=1$ and the input filter
$\mathcal{W}_{2}(s)=\frac{(s+3)^{4}}{\nu(s+50)^{4}}$ (41)
with $\nu$ given in Tab. 1. The inverse of $\mathcal{W}_{2}(s)$ (Fig. 9(a)) serves as an upper bound for the base-controlled system with uncertain pole mass. We then chose $\gamma$ as the inverse of the $\mathcal{H}_{\infty}$ norm of the base-controlled system $\tilde{u}\mapsto\tilde{y}$ with filtering. Based on an empirical study, we observed only minor differences in the results when swapping $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$. Note that the shape of $\mathcal{W}_{2}(s)$ allows for higher gains at moderately high frequencies, which can help to learn a policy that rapidly corrects for displacements of the pole angle. We did not shape $\mathcal{W}_{2}$ at low frequencies, since low-frequency poles require training over very long horizons to capture their behavior.
For comparison, we also trained Residual-LSTM policies — the residual RL architecture (39) with $\mathcal{Q}$ parameterized by a Long-Short Term Memory network (LSTM) [57]. LSTMs are popular recurrent networks known to perform well on RL tasks [58]. We found that Residual-LSTMs were easier to train for this problem than unconstrained Residual-RENs.
7.4.1 Results & discussion
Table 1: Policy hyperparameters.
Policy Architecture Param. $\nu$ in $\mathcal{W}_{2}$ Lip. bound $\gamma$ of $\mathcal{Q}$
Youla-$\gamma$REN $5\times 10^{-4}$ 1.7
Youla-$\gamma$REN (linear) $5\times 10^{-4}$ 1.7
Youla-$\gamma$REN (no filter) - 120
Residual-$\gamma$REN $10^{-2}$ 0.15
Residual-LSTM - -
We first compare the Youla-$\gamma$REN and Residual-LSTM in Figs. 10(a) and 10(b) to investigate the effect of stability guarantees on policy training and long-term test rollouts with model uncertainty. We observe similar qualitative results to Fig. 8. Specifically, the Youla-$\gamma$RENs quickly and smoothly converge to a near-optimal cost during training, and yield stable closed-loop responses over much longer test horizons than the training horizon. In contrast, the Residual-LSTMs often converge to unstable policies during training, causing spikes in their cost curves, and also exhibit closed-loop instability at test time when rolled out over long horizons (e.g., 6 times the training horizon in Fig. 10(b)).
The closed-loop responses under different pole masses (the nominal $m_{p}^{\star}$ and two end points) are depicted in Fig. 11. Both the Youla-$\gamma$REN and Residual-LSTM policies can stabilize the system with a nominal and large pole mass. However, when the pole mass is small, the Residual-LSTM exhibits marginally unstable behavior. That is, the policy rollout of Residual-LSTM is close to that of Youla-$\gamma$REN within the training horizon, but quickly diverges as the time increases. This highlights the importance of model-based stability guarantees in deep RL to achieve reliable long-term test performance.
The sensitivity of the learned policies to variations in the uncertain parameter is shown further in Fig. 10(c). The Residual-LSTM performs well over most of the range, but exhibits instability at small pole masses. The proposed Youla-$\gamma$REN ensures stability and achieves near optimal performance across the whole parameter range. Although its cost is greater than the (optimal) time-varying LQG with known mass, it is very close the time-invariant LQG with known mass, which is reasonable as our policy is also time-invariant. This performance cannot simply be achieved by an LQG controller designed at the nominal mass (shown in gray in Fig. 10(c)) or a Youla-$\gamma$REN controller with linear $\mathcal{Q}$ (see Fig. 12). This suggests that the proposed nonlinear Youla-REN controller has the ability to learn to adapt to different model parameters.
Table 2: Policy architectures for the ablation study (Fig. 12).
Policy Architecture $\mathcal{Q}$ Weight filter
Youla-$\gamma$REN (linear) Youla Linear Yes
Youla-$\gamma$REN (no filter) Youla Nonlinear No
Residual-$\gamma$REN Residual Nonlinear Yes
Figure 11: Policy rollouts on (40) with the same initial state and different pole mass $m_{p}$, where $x_{c}$ is the cart position.
7.4.2 Ablation study
We now conduct an ablation study on the Youla-$\gamma$REN policy class. We trained policies with the three architectures in Tab. 2, all of which have built-in stability guarantees via proper choice of the hyperparameters in Tab 1.
Fig. 12 shows that each of the three Youla-$\gamma$REN components (the Youla parameterization, nonlinear $\mathcal{Q}$, and frequency-weighted filter) is crucial to learn high-performing policies. The linear Youla-$\gamma$REN plateaus at a much higher cost than its nonlinear counterparts due to the poor expressivity of a linear $\mathcal{Q}$. When the weighting filter is removed, the gain of the Youla-$\gamma$REN is restricted equally across all frequencies, leading to a higher cost ceiling. Perhaps most interestingly, there is an extreme drop in performance when switching from the Youla configuration to Residual RL. The main reason is that the gain of $\tilde{u}\mapsto\tilde{y}$ is much smaller than $\tilde{u}\mapsto y$ (Fig. 9), allowing for a more expressive (i.e. larger Lipschitz bound) nonlinear $\mathcal{Q}$ when ensuring closed-loop stability via the small-gain theorem. This reinforces the fact that the policy architecture does matter in deep RL [17]. We learn performant policies with built-in robustness to model uncertainty by combining all three key elements of the Youla-$\gamma$REN policy architecture.
Figure 12: Ablation study on the Youla-$\gamma$REN.
8 Conclusions
We have studied a nonlinear version of the Youla parameterization for nonlinear dynamical systems, and presented a constructive neural realization called the Youla-REN. We have proven that: the parameterization guarantees contracting and Lipschitz closed-loops when the innovations signal is decoupled from the Youla augmentation; if these signals are coupled due to additive disturbances, the disturbed trajectories are constrained within bounded tubes centered on the disturbed-free trajectories; and we cover all contracting and Lipschitz closed loops for certain nonlinear systems. Using RENs for the Youla parameter, empirical studies show that we can learn stabilizing and performant Youla-REN policies on feedback control tasks, even if they are trained with arbitrary (potentially de-stabilizing) cost functions, over short time horizons, and with model uncertainty. For future work, we aim to leverage these advantages and apply Youla-based policy parameterizations to real-world robotic control.
.1 Proof of Theorem 4
.1.1 Claim 1
The result follows by combining Proposition 1 with Theorem 6.
.1.2 Claim 2
We first show that the bound of the innovations signal $\tilde{y}$ (10) is independent of the Youla parameter output $\tilde{u}$, despite there being a feedback loop between $\tilde{u}$ and $\tilde{y}$. Applying observer correctness (A3) to the nonlinear system $\mathcal{G}$, we have
$\begin{split}x_{t+1}&=f(x_{t},\eta_{t},u_{t})+w_{t}\\ &=f_{o}(x_{t},\eta_{t},u_{t},h(x_{t}))+w_{t}\\ &=f_{o}(x_{t},\eta_{t},u_{t},y_{t}-v_{t})+w_{t}\end{split}$ (42)
whose trajectory is denoted by $(x,\eta,u,y,w,v)$. The observer dynamics are
$\hat{x}_{t+1}=f_{o}(\hat{x}_{t},\eta_{t},u_{t},y_{t})$ (9)
with $(\hat{x},\eta,u,y)$ as its trajectory. This shares the same inputs $(\eta,u,y)$ as (42). Since the observer is contracting and Lipschitz (A2), from [59, Thm. 2.8] we have
$\norm{\tilde{x}}_{T}=\norm{x-\hat{x}}_{T}\leq\tilde{\gamma}(\norm{v}_{T}+\norm{w}_{T})+\tilde{\kappa}(x_{0},\hat{x}_{0}),$ (43)
where $\tilde{\gamma}\in\mathbb{R}^{+}$ and $\tilde{\kappa}(x_{0},\hat{x}_{0})\geq 0$ with $\tilde{\kappa}(x_{0},x_{0})=0\ \forall\,x_{0}\in\mathbb{R}^{n_{x}}$. Since the output map $h$ is Lipschitz (A4), we obtain
$\begin{split}\norm{\tilde{y}}_{T}&=\norm{h(x)-h(\hat{x})+v}_{T}\leq\gamma_{h}\norm{\tilde{x}}_{T}+\norm{v}_{T}\\ &\leq\gamma_{0}(\norm{w}_{T}+\norm{v}_{T})+\kappa_{0}(x_{0},\hat{x}_{0})\end{split}$ (44)
where $\gamma_{0}=\tilde{\gamma}\gamma_{h}+1$ and $\kappa_{0}=\gamma_{h}\tilde{\kappa}$ with $\gamma_{h}$ as the Lipschitz constant of $h$.
We now show that the effect of disturbances in constrained within a tube of bounded $\ell_{2}$ norm about the disturbance-free trajectories. Let $\xi_{t}=[x_{t};s_{t}]\in\mathbb{R}^{n_{x}+n_{s}}$ and $q_{t}\in\mathbb{R}^{n_{q}}$ be the states of the base-controlled closed-loop system $(\eta,d,\tilde{u})\mapsto z$ and the Youla parameter $(\eta,\tilde{y})\mapsto\tilde{u}$, respectively. For convenience we denote the state of the total closed-loop system by $\bar{x}_{t}=[\xi_{t};\hat{x}_{t},q_{t}]$. Consider two copies of the system starting at initial states $\bar{x}_{0},\bar{x}_{0}^{\star}$ with inputs $(\eta,w,v)\in\ell$ and $(\eta^{\star},0,0)\in\ell$, respectively, where $x^{\star}_{0}=\hat{x}_{0}^{\star}$. From Assumption (A1) and Lipschitzness of the Youla parameter respectively, we have
$\norm{\Delta z}_{T}$ $\leq\gamma_{1}(\norm{\Delta\tilde{u}}_{T}+\norm{\Delta\eta}_{T}+\norm{w}_{T}+\norm{v}_{T})+\kappa_{1}(\xi_{0},\xi^{\star}_{0}),$
$\norm{\Delta\tilde{u}}_{T}$ $\leq\gamma_{2}(\norm{\Delta\tilde{y}}_{T}+\norm{\Delta\eta}_{T})+\kappa_{2}(q_{0},q^{\star}_{0}),$
for some $\gamma_{1},\gamma_{2}\in\mathbb{R}^{+}$ and $\kappa_{1}(\xi_{0},\xi^{\star}_{0}),\kappa_{2}(q_{0},q^{\star}_{0})\geq 0$ with $\kappa_{i}(a,a)=0$ for $i=1,2$, where $\Delta z=z-z^{\star}$ and similar for all other variables. Combining these inequalities with (44) gives
$\norm{\Delta z}_{T}\leq\gamma(\norm{\Delta\eta}_{T}+\norm{w}_{T}+\norm{v}_{T})+\kappa(\bar{x}_{0},\bar{x}^{\star}_{0})$ (45)
with $\gamma=\gamma_{1}\left(1+\gamma_{2}\max(\gamma_{0},1)\right)$ and $\kappa=\gamma_{1}\gamma_{2}\kappa_{0}+\kappa_{1}+\gamma_{1}\kappa_{2}$. Noting $x^{\star}_{0}=\hat{x}_{0}^{\star}$ implies $\kappa(\bar{x}_{0}^{\star},\bar{x}_{0}^{\star})=0$ completes the proof.
.2 Proof of Theorem 5
.2.1 Useful Result
Consider a contracting system $x_{t+1}=f(x_{t},t)$ with an additive disturbance $\delta(x_{t},t)\in\mathbb{R}^{n}$:
$x_{t}=f(x_{t},t)+\delta(x_{t},t).$ (46)
Let $x,x^{\star}$ be the disturbed and undisturbed trajectories, respectively. The following result is a discrete-time equivalent of [60, Thm. 4]. We omit the proof due to space constraints.
Lemma 1
Suppose $x_{t+1}=f(x_{t},t)$ is contracting with rate $\alpha$ and overshoot $\beta$. If the disturbance satisfies |$\absolutevalue{\delta(x_{t},t)}\leq\bar{\delta}\varepsilon^{t}$ with $\bar{\delta}\in\mathbb{R}^{+}$ and $\varepsilon\in[0,1)$, then we have
[…]
for some $\rho\in[0,1)$ and $C\in\mathbb{R}^{+}$ depending on $\alpha,\varepsilon$.
.2.2 Claim 1
We are interested in the case where $d\equiv 0$ but $x_{0}\neq\hat{x}_{0}$. We first show that the observer error exponentially decays when $d\equiv 0$ regardless of $\tilde{u}$. By setting $w_{t}=0$ and $v_{t}=0$ in (42), the plant dynamics become a copy of the system observer (9) with identical inputs but a different state:
$x_{t+1}=f_{o}(x_{t},\eta_{t},u_{t},y_{t}).$ (48)
[…]
We consider two closed-loop trajectories $(\bar{x}^{a},\eta)$ and $(\bar{x}^{b},\eta)$ with the same input $\eta$ but different initial states $\bar{x}_{0}^{a},\bar{x}_{0}^{b}$. The innovations difference $\Delta\tilde{y}=\tilde{y}^{a}-\tilde{y}^{b}$ decays exponentially as
$|\Delta\tilde{y}_{t}|\leq|\tilde{y}_{t}^{a}|+|\tilde{y}_{t}^{b}|\leq\gamma_{h}(|\tilde{x}_{t}^{a}|+|\tilde{x}_{t}^{b}|)\leq\gamma_{h}\beta_{o}\alpha_{o}^{t}(|\tilde{x}_{0}^{a}|+|\tilde{x}_{0}^{b}|),$
where $\gamma_{h}$ is the Lipschitz constant of the output map $h$. We now estimate the bound of $\Delta\tilde{u}$ via the Youla parameter $\mathcal{Q}$ (11). Note that the trajectories $(q^{a},\eta,\tilde{y}^{a})$ and $(q^{b},\eta,\tilde{y}^{b})$ satisfy
$q_{t+1}^{a}$ $=f_{q}(q_{t}^{a},\eta_{t},\tilde{y}_{t}^{b})+\Delta f_{q_{t}}$
$q_{t+1}^{b}$ $=f_{q}(q_{t}^{b},\eta_{t},\tilde{y}_{t}^{b})$
where $\Delta f_{q_{t}}:=f_{q}(q_{t}^{a},\eta_{t},\tilde{y}_{t}^{a})-f_{q}(q_{t}^{a},\eta_{t},\tilde{y}_{t}^{b})$, and $|\Delta f_{q_{t}}|\leq\gamma_{f_{q}}|\Delta\tilde{y}_{t}|$, since $f_{q}$ has a finite Lipschitz constant $\gamma_{f_{q}}$ w.r.t. $\tilde{y}$ by assumption. Therefore, $\Delta f_{q_{t}}$ decays exponentially, and so by Lemma 1 and the assumption that $\mathcal{Q}$ is contracting with rate $\alpha_{q}\in[0,1)$ and overshoot $\beta_{q}\in\mathbb{R}^{+}$, we have
[…]
where $\max(\alpha_{o},\alpha_{q})\leq\rho_{1}<1$, $C\in\mathbb{R}^{+}$ depending on $\alpha_{o},\alpha_{q}$, and $A_{1}=\beta_{q}\max(1,C\gamma_{f_{q}}\gamma_{h}\beta_{o})$. Since the output mapping $h_{q}$ of $\mathcal{Q}$ is also Lipschitz, we can further conclude that
[…]
where $\tilde{A}_{1}=\gamma_{h_{q}}(A_{1}+\gamma_{h}\beta_{o})$ and $\gamma_{h_{q}}\in\mathbb{R}^{+}$.
We now consider the system mapping $(\eta,\tilde{u})\mapsto z$ under the base controller, which is contracting with some rate $\alpha_{b}\in[0,1)$ and overshoot $\beta_{b}\in\mathbb{R}^{+}$ (A1). By applying Lemma 1 and following a similar procedure as above, we have
[…]
with $\max(\rho_{1},\alpha_{b})\leq\rho_{2}<1$ and $A_{2}\in\mathbb{R}^{+}$ depending on $\tilde{A}_{1},\beta_{b}$ and Lipschitz constants of $f$, $f_{o}$, $f_{b}$, and $k$. Here $\xi_{t}=[x_{t};s_{t}]$ is the state of the base-controlled system $(\eta,\tilde{u})\mapsto z$. By summing (49) with (50) we finally obtain that
$|\Delta\bar{x}_{t}|\leq A\rho_{2}^{t}\left(|\Delta\xi_{0}|+|\Delta q_{0}|+|\tilde{x}_{0}^{a}|+|\tilde{x}_{0}^{b}|\right),$ (51)
where $A=A_{1}+A_{2}$.
.2.3 Claim 2
Now consider the trajectories $(\bar{x}^{a},\eta^{a},z^{a})$ and $(\bar{x}^{b},\eta^{b},z^{b})$. The closed-loop system $(\eta,\tilde{u})\mapsto z$ under the base controller is Lipschitz (A1), as is the Youla parameter $\tilde{u}=\mathcal{Q}(\eta,\tilde{y})$. Therefore, there exist $\gamma_{1},\gamma_{2}\in\mathbb{R}^{+}$ and $\kappa_{1}(\xi^{a}_{0},\xi^{2}b_{0}),\kappa_{2}(q^{a}_{0},q^{b}_{0})\geq 0$ with $\kappa_{i}(a,a)=0$ for $i=1,2$ s.t.
$\norm{\Delta z}_{T}$ $\leq\gamma_{1}\norm{\Delta\eta}_{T}+\gamma_{1}\norm{\Delta\tilde{u}}_{T}+\kappa_{1}(\xi^{a}_{0},\xi^{b}_{0}),$
$\norm{\Delta\tilde{u}}_{T}$ $\leq\gamma_{2}\norm{\Delta\eta}_{T}+\gamma_{2}\norm{\Delta\tilde{y}}_{T}+\kappa_{2}(q^{a}_{0},q^{b}_{0}).$
Since the output mapping $h$ is Lipschitz, we have $\|\Delta\tilde{y}\|_{T}\leq\gamma_{h}\|\Delta\tilde{x}\|_{T}\leq\gamma_{h}(\|\tilde{x}^{a}\|_{T}+\|\tilde{x}^{b}\|_{T})$ and so
$\|\Delta\tilde{y}\|_{T}\leq\frac{\gamma_{h}\beta_{o}(|\tilde{x}_{0}^{a}|+|\tilde{x}_{0}^{b})}{\sqrt{1-\alpha_{o}^{2}}}.$
Combining the above inequalities gives (28) with $\bar{\gamma}=\gamma_{1}(1+\gamma_{2})$, $\tilde{\gamma}=\gamma_{1}\gamma_{2}\gamma_{h}$ and $\kappa=\kappa_{1}+\gamma_{1}\kappa_{2}$.
.3 Proof of Proposition 1
First consider the LTI system (13) with the linear observer from (14). The innovations depend only on the observer error $\tilde{x}$ and external disturbances, since
$\tilde{x}_{t+1}$ $=(A-LC)\tilde{x}_{t}-Lv_{t}+w_{t}$
$\tilde{y}_{t}$ $=C\tilde{x}_{t}+v_{t}.$
This system is contracting and Lipschitz and of the form (29).
Next, consider the fully-observed nonlinear system (17) with innovations (19). We can re-write (19) as
$\psi_{t+1}=w_{t},\quad\tilde{y}_{t}=\psi_{t}$
which is also contracting and Lipschitz and of the form (29).
Finally, consider the partially-observed nonlinear system (1) with the observer (9). Under zero disturbances ($d\equiv 0$, $\tilde{x}_{0}\equiv 0$), by observer correctness (A3) the observer is an exact replica of the plant, hence $\hat{x}_{t}=x_{t}\ \forall\,t\in\mathbb{N}$. Therefore $\tilde{y}_{t}=0\ \forall\,t\in\mathbb{N}$, which is a trivial case of (29).
.4 Proof of Theorem 6
Since all of $\mathcal{T}$, $\mathcal{Q}$, and $\mathcal{G}_{\Delta}$ are contracting and Lipschitz by assumption, then the result follows by the series composition properties of contracting systems and Lipschitz mappings.
.5 Proof of Proposition 2
From the assumptions we have that the incremental dynamics of $\bar{\mathcal{G}}_{\Delta}$ and $\bar{\mathcal{Q}}$ satisfying the following inequalities
where $V_{t}^{g}=V_{t}^{g}(x_{t}^{g},\Delta x_{t}^{g}),V_{t}^{q}=V_{t}^{q}(q_{t},\Delta q_{t})$ are uniformly bounded. Here $x_{t}^{g}=[x_{t};\hat{x}_{t}]\in\mathbb{R}^{2n}$ and $q_{t}\in\mathbb{R}^{q}$ are the internal states of $\mathcal{G}_{\Delta}$ and $\mathcal{Q}$, respectively. By summing the above inequalities and choosing $V_{t}:=V^{g}_{t}+V^{q}_{t}$ as the incremental storage function, we have the closed-loop system is incremental dissipative w.r.t. the following supply rate
$\left[\begin{array}[]{c}\Delta\tilde{y}_{t}\\ \Delta\tilde{u}_{t}\\ \Delta z_{t}\\ \hline\cr\Delta d_{t}\end{array}\right]^{\top}\left[\begin{array}[]{ccc|c}Q_{\bar{y}\bar{y}}+\bar{R}&S_{\bar{u}\bar{y}}^{\top}+\bar{S}&Q_{z\bar{y}}&S_{d\bar{y}}^{\top}\\ S_{\bar{u}\bar{y}}+\bar{S}^{\top}&R_{\bar{u}\bar{u}}+\bar{Q}&S_{\bar{u}z}&R_{d\bar{u}}\\ Q_{z\bar{y}}^{\top}&S_{\bar{u}z}^{\top}&Q_{zz}&S_{dz}^{\top}\\ \hline\cr S_{d\bar{y}}&R_{d\bar{u}}^{\top}&S_{dz}&R_{dd}\end{array}\right]\left[\begin{array}[]{c}\Delta\tilde{y}_{t}\\ \Delta\tilde{u}_{t}\\ \Delta z_{t}\\ \hline\cr\Delta d_{t}\end{array}\right]$
Since the upper-left block of the supply rate matrix is negative definite, by [61, Thm. 2] we can obtain that the closed-loop system is contracting and has a finite Lipschitz bound.
.6 Proof of Theorem 7 & Associated Corollaries
.6.1 Theorem 7.1
Since both of $\tilde{\mathcal{G}}_{\mathcal{K}_{b}}$, $\mathcal{Q}$ are contracting and Lipschitz by assumption, then so is the closed-loop system $\hat{z}=\tilde{\mathcal{G}}_{\mathcal{K}_{b}}(\eta,\tilde{y},\mathcal{Q}(\eta,\tilde{y}))$ by the series composition properties of contracting systems and Lipschitz mappings.
.6.2 Theorem 7.2
The closed-loop system $(\eta,\tilde{y})\mapsto\hat{z}$ consisting of (31) in feedback with $u=\mathcal{K}(\eta,y)$ (16) has dynamics
$\hat{x}_{t+1}$ $=f_{o}(\hat{x}_{t},\eta_{t},h_{\mathcal{K}}(\phi_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t}),h(\hat{x}_{t})+\tilde{y}_{t})$ (52)
$s_{t+1}$ $=f_{b}(s_{t},\eta_{t},h_{\mathcal{K}}(\phi_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t}),h(\hat{x}_{t})+\tilde{y}_{t})$
$\phi_{t+1}$ $=f_{\mathcal{K}}(\phi_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t})$
$\hat{z}_{t}$ $=[\hat{x}_{t};s_{t};h_{\mathcal{K}}(\phi_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t})].$
This system is contracting and Lipschitz by assumption. We will use it to show that $\mathcal{K}$ can be re-written in the form (12) with a contracting and Lipschitz Youla parameter $\mathcal{Q}_{\mathcal{K}}$.
Augment $\mathcal{K}$ with a base controller (8) and observer (9). Using $y_{t}=\tilde{y}_{t}+h(\hat{x}_{t})$ and $\tilde{u}_{t}=u_{t}-\hat{u}_{t}$, the resulting system $\mathcal{Q}_{\mathcal{K}}:(\eta,\tilde{y})\mapsto\tilde{u}$ has the same state dynamics as (52) and is therefore contracting by assumption. Its output map is
$\tilde{u}_{t}=h_{\mathcal{K}}(\phi_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t})-k(s_{t},\eta_{t},h(\hat{x}_{t})+\tilde{y}_{t}),$
hence $\mathcal{Q}_{\mathcal{K}}$ is also Lipschitz since $(\eta,\tilde{y})\mapsto\hat{z}$ is Lipschitz by assumption, $k,h$ are Lipschitz functions (A4), and
$\tilde{u}_{t}=[0,0,I]\hat{z}_{t}-k([0,I,0]\hat{z}_{t},\eta_{t},h([I,0,0]\hat{z}_{t})+\tilde{y}_{t}).$
Moreover, if $\mathcal{Q}=\mathcal{Q}_{\mathcal{K}}$ in the Youla architecture (12), then the control signal $u_{t}=\hat{u}_{t}+h_{\mathcal{K}}(\phi_{t},\eta_{t},y_{t})-\hat{u}_{t}$ is unchanged from $u=\mathcal{K}(\eta,y)$. Hence $\mathcal{K}$ can be written in the form (12) with $\mathcal{Q}=\mathcal{Q}_{\mathcal{K}}$ contracting and Lipschitz.
.6.3 Corollary 1
The closed-loop system $(\eta,w)\mapsto z$ consisting of (17) in feedback with $u=\mathcal{K}(\eta,y)$ (16) has dynamics
$x_{t+1}$ $=f(x_{t},\eta_{t},h_{\mathcal{K}}(\phi_{t},\eta_{t},x_{t}))+w_{t}$ (53)
$\phi_{t+1}$ $=f_{\mathcal{K}}(\phi_{t},\eta_{t},x_{t})$
$z_{t}$ $=[x_{t};h_{\mathcal{K}}(\phi_{t},\eta_{t},x_{t})].$
This system is contracting and Lipschitz by assumption. Make the substitutions $x_{t}=\hat{x}_{t}+w_{t-1}$ and $\tilde{y}_{t}=w_{t-1}$ which follow by (10), (19) since $y_{t}=x_{t}$. Then (53) becomes
$\hat{x}_{t+1}$ $=f(\hat{x}_{t}+\tilde{y}_{t},\eta_{t},h_{\mathcal{K}}(\phi_{t},\eta_{t},\hat{x}_{t}+\tilde{y}_{t}))$ (54)
$\phi_{t+1}$ $=f_{\mathcal{K}}(\phi_{t},\eta_{t},\hat{x}_{t}+\tilde{y}_{t})$
$z_{t}$ $=[\hat{x}_{t}+\tilde{y}_{t};h_{\mathcal{K}}(\phi_{t},\eta_{t},\hat{x}_{t}+\tilde{y}_{t})].$
This system is exactly (52) with the observer $\hat{x}_{t}=f(y_{t},\eta_{t},u_{t})$ but for the addition of $\tilde{y}_{t}$ in the output $z_{t}$, which does not change the fact that it is contracting or Lipschitz. The result then follows by Theorem 7.
.6.4 Corollary 2
The closed-loop system $(\eta,d)\mapsto z$ consisting of (1) in feedback with $u=\mathcal{K}(\eta,y)$ (16) has dynamics
$x_{t+1}$ $=f(x_{t},\eta_{t},h_{\mathcal{K}}(\phi_{t},\eta_{t},h(x_{t})+v_{t}))+w_{t}$ (55)
$\phi_{t+1}$ $=f_{\mathcal{K}}(\phi_{t},\eta_{t},h(x_{t})+v_{t})$
$z_{t}$ $=[x_{t};h_{\mathcal{K}}(\phi_{t},\eta_{t},h(x_{t})+v_{t})].$
This system is contracting and Lipschitz by assumption. Take (55), re-label $x$ as $\hat{x}$, and choose $v_{t}=\tilde{y}_{t}$, $w_{t}=L(\eta_{t},\tilde{y}_{t})$. Then the dynamics are exactly (52) when the base controller and observer have the form (32). The above transformation is Lipschitz (because $L$ is Lipschitz), hence if (55) is contracting and Lipschitz so too is (52). The result follows by Theorem 7.